Shift report 20010925 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Sep 25 2001 - 15:59:14 EDT

  • Next message: Flemming Videbaek: "new shift list"

    ------------------------------------------------------------
    Date: Tue Sep 25 15:59:13 2001, Shift: 08:00-16:00
    Supervisor: Eun-Joo Kim, with: FV, DB
    Log book#9 pages 39 - 43
    
    Shift Summary:
    ------------------------------------------------------------
    1. 2 hours 45 min beam hour
    
    2. Data Run : 4927   20k trigger 1
                  4932   21k trigger 1
                  4933  128k trigger 1   -> 169k MRS(90) FS(3) 
                  4940   now runnig
    
       Pedestal Run : 4930
       TPC Pedestal Run : 4929
       DVM Run : 4928, 4931
    
       Run for Background study : 4934 FS  5-deg 
                                  4935 FS  7-deg
                                  4936 FS  9-deg
                                  4937 FS 11-deg  
                                  4939 FS  4-deg
    3. T5 Anode 1 : now OK
                    Tip??? -> please check T5 low voltage
                              please check T5 trip or not
       
       FV updated BRAHVO for HV trip problem.   
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Sep 25 2001 - 15:59:44 EDT